News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 25 July 2003 9.30am EDT

During the last week the observing schedule was replanned to include an observation of the the X-ray Flash source H2777 that was accepted as a 3 day turn-around Target Of Opportunity (TOO) on 23 July. Loads were replanned and uplinked on Jul 24 for the observation on Jul 25. A portion of an observation of M8 was impacted by the TOO and has been rescheduled for next week.

As previously reported, the swap from Inertial Reference Unit-1 (IRU-1) to IRU-2 was completed as planned on July 19 and the mission is proceeding nominally under IRU-2 control. IRU-1 is now configured as the backup IRU for use in the event of a safemode transition. A set of maneuvers was performed following the swap to verify the updated IRU calibration and completed with acceptable results. The swap was performed during a perigee passage and the science program was restarted following radiation belt exit.

A real-time procedure was completed on July 22 to perform a manual unload of momentum. The unload was performed at an attitude that ensured a cooler starting temperature than prior anomalous unloads. Thruster performance was nominal during the unload.

The observing schedule for next week is shown below.

------------------------------------------
Radiation Belts                     Jul 27
M8                     ACIS-I
ZWCL1215               ACIS-I       Jul 28
J1314+3649             ACIS-S
NGC5678                ACIS-S
PC 1447+4750           ACIS-S
HR 6406                HRC-I
V404 CYG               ACIS-S
NGC7465                ACIS-S       Jul 29
3810                   ACIS-S
M33 X-8 (Nucleus)      HRC-S
Radiation Belts
H1743-322              ACIS-S/HETG  Jul 30
A3571                  ACIS-S       Jul 31
NGC 4168               ACIS-S
EXO 1747-214           ACIS-S
OGLE-1999-BUL-32       ACIS-S
G11.2-0.3              ACIS-S       Aug 1
AX J1845-0258          ACIS-S
Radiation Belts
SDSS J0654.11-001533.4 ACIS-S       Aug 2
MS0440.5+0204          ACIS-S
NGC 988                ACIS-S
A209                   ACIS-I       Aug 3
F01004-2237            ACIS-S
SGR 1806-20            ACIS-S

------------------------------------------

All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports